OpenStudy (anonymous):

solve for x: ln(8x+1) = 2ln(x+2)

OpenStudy (saifoo.khan):

ln(8x+1) = ln(x+2)^2 cancel the lns 8x+1 = (x+2)^2

OpenStudy (saifoo.khan):

Solve..

OpenStudy (anonymous):

oh wow! its that simple?!

jimthompson5910 (jim_thompson5910):

remember to check all possible solutions

OpenStudy (anonymous):

so x=1 and x=3?